OpenStudy (anonymous):

Please help Pre-Calc 1.Write the expression as the sine, cosine, or tangent of an angle. sin 8x cos x - cos 8x sin x 2.Rewrite with only sin x and cos x. cos 2x - sin x

OpenStudy (anonymous):

looks like the "subtraction angle formula" for sine

OpenStudy (anonymous):

so how would I do it? Can you please show me the step by step process of how to do it?

OpenStudy (anonymous):

\[\sin(A-B)=\sin(A)\cos(B)-\cos(A)\sin(B)\] put \(A=8x,B=x\) and go from the right side to the left side

OpenStudy (anonymous):

actually there is not much to do you get \[\sin(8x-x)=\sin(7x)\] and you are done

OpenStudy (anonymous):

So would the second question be cos2x + sin2x +sinx?